VISTA IOT MEETING 2010-08-24 ============================ Minutes by: M. Rejkuba ====================== August 24, 2010 09:00 Chile, 15:00 Garching, 14:00 UK ---------------------------------------------------------- CONTENTS: 0: Presence at the meeting 1: VISTA status and planned activities; pending visit to Chile of G. Hussain et al. 2: Status of various software 2.1. SADT 2.2. OT 3: QC issues, status of the reprocessing (WHU, CASU) 4. Progress of the surveys (MRE) 5: Action Items review 6: Other issues 7: Date of the next meeting =========== === 0. PRESENCE ======================================================== =========== Santiago - S. Mieske ................................... [SMI] Garching - M. Rejkuba ................................... [MRE] W. Hummel ................................... [WHU] M. Arnaboldi ................................... [MAR] via Phone - J. Emerson ................................... [JEM] CASU: M. Irwin ................................... [MIR] J. Lewis ................................... [JLE] S. Hodgkin .................................. [SHO] ================================================================= === 1. VISTA status and planned activities ======== ================================================================= SMI reports on the VISTA status and planned activities: 1.1 Camera recentering The camera has been recentered, which fixed the problem with the focus. The image quality looks better and TSZ circulated recently an e-mail with the results. After centering the camera in the v blocks the focus gradient seems to be gone. 1.2 Calibration units The calibration units for dome flats are operational and calibration database should be updated. WHU asks what it means that the calibration untis are operational. Does that mean that there is no need to go to the dome? Also what about the stability? The lamp needed to be switched on for at least 1h to stabilize. There were some recent flats in H and K that were saturated. With DIT=1sec there were 15000 counts, but when the image is read out this is actually double and then saturated. SMI states that the lamp can be adjusted, but he will need to verify with TSZ about the question of stability. JEM mentions that different transformers were used to stabilize the lamp and change the levels as needed. Also the old lamp needed some time to stabilize. New AI for SMI: verify with TSZ about the flatfield lamp stability and adjustment for the correct counts for H and K flats. 1.3 VISTA Operations (SciOps) SMI reports that VISTA is currently operated by TIOs. The astronomers in the survey operations group are only T. Szeifert, V. Ivanov and S. Mieske (TSZ, VIV, SMI). In principle the shift coordinator should take care of the daily tasks, but the idea is now to increase the survey operations group in order to have more people that could be on the mountain in order to take better care of the follow-up of VISTA. MRE asked who the new people are. The new members of the survey operations group will be: Henry Boffin, Silvain Guieu, Giovanni Carraro. They will get full training (only few nights) and will be mostly in charge of VISTA during day time. In addition shift coordinators remain in charge (responsible) for the follow-up of daily tasks for VISTA (check tickets, vista e-mail account, etc.). However, as of September they should train more people, so that they can delegate some of the actual work to day astronomer. 1.4. Overheads The overheads for the VVV OBs execution are significantly larger than estimated by the P2PP. TSZ discussed this with the operatiors and it turned out that very often they do a manual AO correction at the beginning of each OB. They were asked not to do that. This however, may result in poorer image quality. MRE commented that if the AO correction is necessary (these are OBs with too short (DIT+2)*NDIT integration time to have AO running in parallel), then the AO priority should be set to HIGH and the appropriate additional overhead charged that accounts for this necessary AO correction. 1.5 visit of G. Hussain et al. to Chile Regarding the part of the agenda that mentioned to discuss "Pending visit of G. Hussain et al. to Chile" - due to the fact that this visit should be coordinated between V. Ivanov and G. Hussain, and neither of the two could attend the meeting, it was decided that this should be done via direct e-mail exchange and not at the IOT meeting. 1.6. Mirror coating degradation JEM points out that the zero points have degraded strongly. The trend over the last 9 months can be seen on the CASU technical web page: http://casu.ast.cam.ac.uk/surveys-projects/vista/technical/photometric-properties CASU web page shows that ZPs went down by about 20% in Z, but only few % for K. The bluer the band the worse the degradation. This trend is very different from the 11 months trend that the previous coating had. That may point in the direction of someting to do with the new coating. The difference with the previous one is in the CO2 cleaning. JEM points out that this needs to be raised at the next technical telecom. It is important to understand why there is such a fast degradation and then to decide on the next re-coating. SMI asks for the timescale for the recoating. The reply is that this is driven by the scientific requirements. The specifications were done for the IR, but the worst degradation is for bluer bands. Nevertheless the users are getting shallower data and this will have impact on the science. Of course also the re-coating will have some impact, because the telescope will not be available for operations for at least 1 week. According to SMI the re-coating frequency for UTs is ~18 months. It is not clear how frequently it was planned for VISTA. SMI asks if Garching QC has indications of similar degradation. WHU answers that the trends for the ZPs are not checked regularly. JEM mentions that the secondary was never recoated. MRE asks whether the decrease in sensitivity comes from the primary or secondary. JEM replies that at the technical maintenance telecom there was some mention about the degradation of the primary based on visual inspection. However, it is true that it is at the moment not known whether the degradation comes (only) from the primary. This may need to be checked. New AI for SMI to discuss with TSZ and engineers in order to investigate and possibly schedule the next re-coating. ======================================= ============== 2. Software tools status ===================== ======================================= 2.1. SADT JEM has not yet delivered the final version of SADT. He reports on the latest changes and says that he is currently trnasfering all the changes to CVS. This should be done by the next week. After that he will deliver the tool to M. Hilker as agreed. One thing that however has not yet been checked is that the latest changes have not impacted on the VST module. This check will be done only after the tool is in the CVS. SADT seems to be much faster now and therefore it is more difficult to identify what went wrong if there is a problem detected. MRE reports on the problem that VIKING OBs had due to scratches on the plates used for the GSC2 catalogue. Due to that no AO and guide stars were available. The OBs had to be revised. This was done using alternative catalogue (UCAC3) and this was also used then to do some comparison between the catalogues. JEM reminds that there is now a possibility to set the correct magnitude limits in the config file. The current limits were tuned for GSC2 but may not be appropriate for the other catalogues and this will need to be tested. 2.2. OT & P2PP: MRE reports on the latest changes and development of the OT3. The P2PP version 3.1.4, including the twilight constraint was used for the preparation of the OBs for P86. Therefore the new version of the OT will need to be issued before October to use this constraint. In addition there were some discussions and progress made on improving the reporting and in particular the OB history and events connected with execution and subsequent QC grade and status modifications for the OBs. This still needs to be implemented and should be fixed in the next OT version to be delivered by the end of September. ===================================================== ======= 3. QC issues, data products issues ============ ====================================================== 3.1 Radioactive events & persistence WHU analyzed radioactive events. An e-mail with the results was circulated to the IOT members and the QC web page reports on the findings. The persistence report was circulated to the IOT by WHU. This should however still be added to the web pages. 3.2 Images with high ellipticity WHU expressed his concerns about the large number of OBs with ellipticity close to 0.2. Currently the agreement is that if there are 4 or more detectors with ellipticity >0.2, then this should be repeated. However, there are many OBs that have only 2-3 detectors with e>0.2 and all the other detectors have ellipticity that is just below 0.2 threshold. Worry is that the PIs may come back 1/2 year later and say that these data are not good enough for the scientific goals. SMI suggests that perhaps it may be necessary to implement QC criteria connected with the mean ellipticity. He asks whether these cases where most detectors had ellipticity just below 0.2 limit were done in particularly good seeing. WHU answers no - these were taken on an average night, not under the best seeing conditions. MRE asks whether PIs receive ellipticity measurements. They receive reports from CASU and it is interesting to know whether among the parameters they get ellipticity of the stellar sources. In addition, there was an e-mail circlulated to all the PIs asking them to eye-ball the data and to send information if some data/images need to be deprecated. Question is how many PIs contacted CASU about deprecation of data. They did not contact ESO so far... JLE confirms that the ellipticity is one of the parameters that is given to the PIs. So far none of the PIs replied about the deprecation of images. MRE will contact the PIs to remind them that they are responsible to verify the quality of the data, becasue ESO does only spot checks. 3.3 Recent twilight flats WHU reports that some recent twilight flats needed to be discarded due to eitehr too low counts or non-continuous decay that is reminiscent of the flats taken with thin cirrus-es. There were many flats with irregular decrease. SMI will re-iterate to telescope opeartors not to take sky flats unless it is clear. The operations rule is not to take flats and not to take standard star observations if there are cirruses. However, apparently this was not always very strictly adherred to. WHU mentions that for some nights when these discontinuous flat sequences were found ASM showed clear flux variations (indicative of clouds), but on some nights not. Possibility is that in the direction where ASM measures there were no clouds, but in the direction of the twilight field of VISTA there were some/few thin cirruses that went unnoticed. 3.4 Short term variability of read-out registers WHU reports on the finding that the vertical stripes (coming from the 16 readout registers) appear to vary also on 2-3h time-scale. When he reduced night sky flat with twilight flat the gain variations were visible. MIR comments that the sky subtraction should take out the small gain variations. ===================================================== ======= 4. Survey progress ============ ====================================================== MRE gave a quick overview of the status of various surveys. Conclusion is that due to bad weather and some technical issues, there is still a lot to do and there does not seem to be danger of empty queues (for decent weather conditions). The survey completion has fallen back with respect to prediction. Here are some numbers (added after the IOT meeting): 1. 179.B-2002(B), VVV OBs: 1040 completed / 2460 submitted - completion rate 42% 142 OBs have grade B 1102 OBs attempted (status: C, M, A) time: 300h allocation; completed 181.8h - completion rate 61% 2. 179.B-2003(C) VMC OBs: 20 completed / 39 submitted - completion rate 51% 4 OBs have grade B 21 OBs attempted time: allocation 30h, 17h completed - completion rate 57% 3. 179.A-2004(B), VIKING OBs: 154 completed / 268 submitted - completion rate 57% 29 OBs have grade B 161 OBs attempted time: 320h allocation; 175.5h completed - completion rate 55% 4. 179.A-2005(B), Ultra-VISTA OBs: 139 completed / 141 submitted - completion rate 99% 27 OBs have grade B 140 OBs attempted time: 160h allocation; 157.8h completed - completion rate 99% 5. 179.A-2006(B), VIDEO OBs: 40 completed / 100 submitted - completion rate 40% 4 OBs have grade B 40 OBs attempted time: 100h allocation; 40.1h completed - completion rate 40% 6. 179.A-2010(B), VHS OBs: 966 completed / 1088 submitted - completion rate 89% 157 OBs have grade B 986 OBs attempted time: 410h allocation; 325.5h completed - completion rate 79% Ultra-VISTA is not observable any more. It is almost completed. VHS is also mostly done, but there are still more than 100 OBs (~85h) that can be used for the nights with somewhat worse seeing. Most of the observations are now concentrated on completing as much as possible VVV, and VIKING, which are lagging behind the predicted completion. VIDEO is the least completed, but its observing season (visibility) period just started. ===================================================== ======= 6. Action Items review ============ ====================================================== OLD AIs: -------- AI 2009-11: TSz should find out if the constraints sets could be carried with the OBs, to the OS and the fits header of the VIRCAM files. Otherwise it is not easy to propagate these data by interfaces. 2009-07-29: TSz iterating with R. Schmutzer, but DFI (T. Bierwirth) will also need to be involved. For surveys, given the small range of programs it is expected that there are not too many different OBs, and this might need to be followed in a different/manual way in the beginning - this could imply a major change for the tools that cannot be done in the last moment. 2010-03-23: the keywords have been implemented except for the moon; there is a ticket with details; some testing done by VDI, no errors found; keep ongoing, try to complete during my next shift end of Mar - early Apr. 2010-04-23: VDI reports one formula is missing 2010-05-27: Corresponding PPRS is still pending. VIV asks to keep this AI open until the PPRS is closed. 2010-06-29: No changes with respect to last meeting. VIV asks to keep this AI kept open/pending until the PPRS is closed. 2010-08-24: VIV not present at the meeting. It is not known if there was any progress. Therefore decided to keep this pending while waiting for final status from VIV. STATUS: PENDING ** AI 2009-14: GHu to check that the seeing distribution function in the scheduler for VIRCAM takes into account the finite (0.51") nominal resolution of the instrument. 2010-03-23: GHu reported that it is not clear if TaToo will be used for scheduling time on VISTA; decided to keep this AI open until it becomes clear how VISTA time will be scheduled. 2010-04-23 no news 2010-05-27: TSZ states that for the scheduling the seeing distribution is _not_ considered in the requested band. MRE notes that there may be some chance that this is updated by ESO. 2010-06-29: GHU reports that Gautier Mathys is reluctant to make these changes if the open time is scheduled in the specific time slots. We need to formalize and few weeks the most efficient way to do scheduling. For the surveys this is with the VIV simulation rather than with TaToo. VIV states that he will be on Paranal 2 weeks. He points out that it is either necessary to re-create the database for the next semester based on SMPs or to wait for the PIs to submit the OBs. GHU says that we cannot wait for the PIs and this should be discussed. Gautier needs to be involved in this decision as well. Best is to discussed by phone next week. Decision on how to do scheduling for surveys + open time has to be taken before the next IOT. VIV points out that he would like to have it done before going down on July 10. 2010-08-24: Since the last IOT there is an agreement that the simulation tool of VIV will be used for the scheduling of surveys. Since this tool has implemented the proper ranking and includes the instrument resolution in the seeing distribution this AI becomes obsolete. Thus closed. STATUS: CLOSED ** AI 2010-06: TSz+VDI to improve the TIO training, to prepare an operational manual 2010-05-27: Operations wiki has been updated. More training needed. 2010-06-29: Not much was done due to problems with IQ + intervention + work on IP. Keep the AI open. 2010-08-24: SMI comments that the training is an ongoing task. Therefore it seems that this could be closed. It is not clear about the status of the operational manual mentioned in the AI. Thus for the moment kept as PENDING. Comment from VIV and TSZ on the status required. STATUS: PENDING ** AI 2010-09: photometric zeropoints as calculated by the pipeline. 2010-05-27: Ongoing, see corresponding discussion this meeting 2010-06-29: TSZ points out that this is not well defined AI. MIR adds that the currently existing plots already would have shown the problems with zero points from pawprints to tiles. It is not clear whether this is about the zero points from Garching pipeline? WHU states that Garching pipeline has de-blending on. MIR: There is no significant increase in the scatter with respect to what one would expect from the random noise. It is agreed in the end to send one example data set and then the comparison should be done directly between Garching and CASU. 2010-08-24: This is still pending. WHU should send some example to JLE/MIR in order to do the checks and clarify what the problem actually is about. STATUS: PENDING ** AI 2010-11: WHU to distribute the persistence report to the whole IOT. 2010-08-24: DONE STATUS: CLOSED ** AI 2010-12: VIV put in the User manual persistence amplitude and slope. Add to the instrument page the persistence report. 2010-08-24: MRE reports that the persistence is mentioned in the latest version of the manual, but not the amplitude and slope. The persistence report still needs to be added to the instrument web page (link in the news?). STATUS: PENDING ** NEW AIs: -------- AI 2010-13: SMI to verify with TSZ the flat field lamp stability and the adjustment of the correct voltage for H and Ks flats. AI 2010-14: SMI to raise the issue of degradation of coating (based on degradation of photometric zero points). The next coating should be scheduled. AI 2010-15: MRE to contact the survey PIs and remind them that the QC checks on the science data need to be done by the survey teams and any issues should be reported to CASU and usd-help@eso.org. ======================================= ============== 7. Other issues ===================== ======================================= WHU asks when the first data products are to be expected from the PIs. MAR replies that the current schedule forsees the Phase 3 in March 2011, i.e. about 1 year after the official start of the surveys. ======================================= ============== 8. Date of the next meeting ===================== ======================================= TUESDAY, 5. October, 2010 9:00 Chile, 15:00 Garching